OpenStudy (kitkat1):

Medal for Best answer! Which of the following exponential functions goes through the points (1, 6) and (2, 12)? f(x) = 3(2)x f(x) = 2(3)x f(x) = 3(2)−x f(x) = 2(3)−x

OpenStudy (kitkat1):

Can someone plz help me

OpenStudy (anonymous):

What do you think? :D

OpenStudy (kitkat1):

The last one

OpenStudy (kitkat1):

But not really sure

OpenStudy (anonymous):

ya, it is :D

OpenStudy (kitkat1):

are u sure why?

OpenStudy (anonymous):

definitely not D

OpenStudy (kitkat1):

Then B?

OpenStudy (anonymous):

lets try \[f(x) = 3(2)^x\]

OpenStudy (anonymous):

and then we can try \[f(x)=2(3)^x\] and see which one works, instead of guessing A, B, ...

OpenStudy (anonymous):

both give you \[f(1)=6\] the question is, which one gives \(f(2)=12\)

OpenStudy (anonymous):

A would be \[f(2)=3\times 2^2\] and B would be \[f(2)=2\times 3^2\] which one equals 12?

OpenStudy (kitkat1):

A

OpenStudy (kitkat1):

So the answer is A?

OpenStudy (kitkat1):

satelite 73